If each of the circles has a diameter of 12 meters, what is the area of the region inside the four circles? (Use π = 3.14. Round your answer to one decimal place.)

Answer by lwsshak3(11628)

If each of the circles has a diameter of 12 meters, what is the area of the region inside the four circles? (Use π = 3.14. Round your answer to one decimal place.)
***
Area of a circle=πd^2/4, d=diameter
Area of 4 given circles with diameter of 12 m=4*3.14*12^2/4= 452.2 m^2